OpenStudy (anonymous):

A balanced die is rolled twice. Are the events "the sum of the two rollsis 8" and "the first roll comes up 3" independent?

OpenStudy (anonymous):

no they are not, but this will take some work

OpenStudy (anonymous):

the probability that the sum is and 8 is \(\frac{5}{36}\) by counting the number of ways to get an 8 on two dice

OpenStudy (anonymous):

the probability that the sum is an 8 given that the first die is a 3, here we also count. the sample space now has 6 element namely (3,1) (3,2) (3,3) (3,4) (3,5) (3,6) of those 6, one has a total of 8 so the probability that you roll and 8 GIVEN that the first die is a 3 is \(\frac{1}{6}\)

OpenStudy (anonymous):

now since \(\frac{5}{36}\neq \frac{1}{6}\) that means the events are DEPENDENT

OpenStudy (anonymous):

i.e. not independent

OpenStudy (anonymous):

events A, B are independent means \(P(A|B)=P(A)\) and since this is not the case here, they are dependent

OpenStudy (anonymous):

Thanks a lot!!!!

OpenStudy (anonymous):

yw